OpenStudy (anonymous):

If a2=-4 and a5=-3 find the first and the sixth term of this arithmetic sequence

OpenStudy (amistre64):

an = a1 + d(n-1) you can shift the indexes to use these to find the sequence with

OpenStudy (amistre64):

a2 =-4 a5=-3 a1 a4 an = a1 + d(n-1) a4 = a1 +d(4-1) -3 = -4 +d(3) solve for d and reuse the setup